# Theory and Experimental Demonstration of Wigner Tomography of Unknown Unitary Quantum Gates

### v1

@@ IX Conclusion and Discussion
The tomography approach presented here consists of two parts: experimental acquisition and classical post-processing of data. The experimental part provides partial information (i.e., droplets corresponding to scaled processes) about the experimental process. In the post-processing stage, the information contained in these experimental scaled process droplets is combined to reconstruct the unknown process. We introduced a reconstruction algorithm for a single-qubit system and demonstrated its potential to rebuild the experimental process matrix from noisy droplets with a high signal-to-noise ratio (see Sec. V). Additionally, we showed how the reconstruction algorithm can be integrated with an optimization routine, highlighting the benefits gained from this combined approach. The reconstruction algorithm can also be extended to an $N$ qubit system, as discussed in Sec. VIII.
@@
The Wigner tomography approach for unknown processes fills a gap in the general theory work of Wigner tomography but is resource-intensive in comparison to the conventional tomography approaches [15] both in terms of the number of qubits and the number of experiments. In general, for Wigner tomography of an $N$ qubit unknown process, $12N+$ 1 qubits are required. The total number of experiments in terms of the number of controlled-$\mathcal{G}$ operations can be significantly cut down by employing an adaptive approach discussed in Sec. VII. In the near term, for devices with a limited number of qubits, where the scalable error correction is not viable, the long experimental implementation time of the CSWAP gates used in this approach also results in a loss of signal-to-noise ratio of the tomographed droplets due to decoherence effects and experimental imperfections.
@@
The unknown process tomography approach presented herein serves as a tool for analyzing a “black box” containing one or more unknown gates and whose inner workings are hidden. In one scenario, a quantum computer is considered that is capable of executing all required gates for the tomography with high fidelity. In this case, the black box can be plugged directly into the tomography circuits to perform the complete process tomography to extract full information about the black box. In a second scenario, consider a case where some partial information about the black box may be available. Here, the adaptive version of the tomography approach (see Sec. VII) can be employed to extract full information about the black box, where such partial information can be leveraged to design the $\text{c}\mathcal{G}$ rotation to efficiently estimate the actual process with fewer experiments. Finally, in a third scenario, consider a case where the target process is fully known and where the tomography approach is used for a “quality check”, i.e., to assess the experimental performance of a given target gate $U_{\text{tar}}$ on a particular quantum device. In this case, only one tomography experiment (with $\text{c}\mathcal{G}=\text{c}U_{\text{tar}}$, see Fig. 10) is sufficient to estimate the performance of the actual gate, i.e., experimental implementation of the target gate.
@@ D.1 Correlation matrix
where $U$ is a reconstructed process matrix and $U_{t}$ is a target process matrix, which can be constructed based on Eq. 20.

### v2

@@ IX Conclusion and Discussion
The tomography approach presented here consists of two parts: experimental acquisition and classical post-processing of data. The experimental part provides partial information (i.e., droplets corresponding to scaled processes) about the experimental process. In the post-processing stage, the information contained in these experimental scaled process droplets is combined to reconstruct the unknown process. We introduced a reconstruction algorithm for a single-qubit system and demonstrated its potential to rebuild the experimental process matrix with a high signal-to-noise ratio from noisy droplets (see Sec. V). Additionally, we showed how the reconstruction algorithm can be integrated with an optimization routine, highlighting the benefits gained from this combined approach. The reconstruction algorithm can also be extended to an $N$ qubit system, as discussed in Sec. VIII.
@@
The Wigner tomography approach for unknown processes fills a gap in the general theory work of Wigner tomography but is resource-intensive in comparison to the conventional tomography approaches [15] both in terms of the number of qubits and the number of experiments. In general, for Wigner tomography of an $N$ qubit unknown process, $12N+$ 1 qubits are required. The total number of experiments in terms of the number of controlled-$\mathcal{G}$ operations can be significantly cut down by employing an adaptive approach discussed in Sec. VII. In the near term, for devices with a limited number of qubits, where a scalable error correction is not viable, the long experimental implementation time of the CSWAP gates used in this approach also results in a loss of signal-to-noise ratio of the tomographed droplets due to decoherence effects and experimental imperfections.
The unknown process tomography approach presented herein serves as a tool for analyzing a “black box” containing one or more unknown gates whose inner workings are hidden. Here, we consider three different scenarios:
  • In a first scenario, a quantum computer is considered that is capable of executing all required gates for the tomography with high fidelity. In this case, the black box can be plugged directly into the tomography circuits to perform the complete process tomography to extract full information about the black box.
@@
  • In a second scenario, consider a case where some partial information about the black box may be available. Here, the adaptive version of the tomography approach (see Sec. VII) can be employed to extract full information about the black box, where such partial information can be leveraged to design the $\text{c}\mathcal{G}$ rotation to efficiently estimate the actual process with fewer experiments.
  • In a third scenario, consider a case where the target process is fully known and where the tomography approach is used for a “quality check”, i.e., to assess the experimental performance of a given target gate $U_{\text{tar}}$ on a particular quantum device assuming that the error is small. In this case, only one tomography experiment (with $\text{c}\mathcal{G}=\text{c}U_{\text{tar}}$, see Fig. 10) is sufficient to estimate the performance of the actual gate, i.e., experimental implementation of the target gate.
[…]
The total number of shots is defined as $N_{tot}=N_{p}\cdot N_{s}$, where $N_{p}$ is the number of experiments required for a given tomography method, and $N_{s}$ is the number of shots per experiment. For example, the standard tomography approach [15] requires $N_{p}=12$, whereas complete Wigner tomography (first scenario) requires $N_{p}=208$. The latter value is based on a Lebedev grid with $n=26$, requiring $2$ circuits for process tomography (considering four circuits to prepare maximally mixed states with shots evenly distributed, see Fig.10), repeated $4$ times for different rotations $\mathcal{G}$. Differences in $N_{p}$ across methods are compensated by adjusting $N_{s}$ to maintain the same $N_{tot}$.
Figure 18: Comparison of different versions of the Wigner tomography method (see text) and the standard process tomography approach. The corresponding plot with standard deviation is shown in Fig. 28 of Appendix G.
For the adaptive approach (second scenario), Fig. 18 shows the results after two iterations. In the first iteration, the rotation axis of $U_{\text{gue}}$ is assumed to be tilted by $30^{\circ}$ away from the rotation axis of the target operator $U_{\text{tar}}$ (whereas the rotation angles of $U_{\text{gue}}$ and $U_{\text{tar}}$ are assumed to be identical). The operator $\mathcal{G}$ is chosen to be identical to $U_{\text{gue}}$. The total number of shots, $N_{tot}$, was evenly split between two iterations, with $N_{tot}/2$ allocated to each. This ensures that the data points presented in Fig. 18 effectively correspond to the same total number of shots across all approaches.
For the non-iterative approach (third scenario), only one experiment with rotation $\mathcal{G}=U_{\text{tar}}$ was performed. In this simulation, the rotation axis of $U_{a}$ is tilted away from the rotation axis of $U_{\text{tar}}$ only by a small angle of $2^{\circ}$ (while the rotation angle of the actual operator $U_{a}$ and of the target operator $U_{\text{tar}}$ is assumed to be identical). In the case of the standard process tomography approach, Qiskit and Forest-Benchmarking functions were used to reconstruct the channel matrix and to compute the unitary closest to a quantum process specified by a channel matrix [30]. For all the approaches, the process fidelity was computed using the definition shown in Eq. (58).
The results show that the non-iterative Wigner tomography method, where only a small error between the target operator $U_{\text{tar}}$ and the actual operator $U_{\text{a}}$ is assumed (third scenario), performs better than other Wigner tomography approaches while requiring one-fourth the number of experiments $N_{p}$ compared to the complete Wigner tomography approach (first scenario) for the same $N_{tot}$. However, the standard tomography approach achieves higher precision than Wigner tomography methods for the same $N_{tot}$. The plot in Fig. 18 is replotted in Fig. 29 of Appendix G with a reduction of the number of shots $N_{s}$ for the standard approach by a factor of 4 (purple dashed line) and by a factor of 8 (green dashed line). This illustrates that the standard approach only requires between four and eight times less total number of shots to reach a comparable precision as the non-iterative Wigner tomography method (blue solid line). This difference possibely arises because, for a single-qubit process tomography, Wigner tomography requires measuring two-qubit expectation values (see Eq. (27)), while the standard approach only needs one-qubit expectation values [15].
X Code availability
The Matlab and Python codes used for the tomography approach are available in the GitHub repository [31]. This repository also contains the code for plotting the spherical droplets from the Lebedev grid and for performing standard process tomography. The Python code uses Qiskit [32], an open-source development for simulations and experiments that is directly usable on IBM quantum computers and can be modified for other quantum computing platforms.
@@ D.1 Correlation matrix
where $U$ is a reconstructed process matrix and $U_{t}$ is a target process matrix, which can be constructed based on Eq. (20).
@@ F.3 $\mathrm{Z}$ gate
Appendix G Additional plots for the comparison of Wigner and standard process tomography approach
Figure 28: Replot of Fig. 18 including standard deviations, with vertical bars representing the standard deviation for each data point.
Figure 29: Replot of Fig. 18 with fewer shots $N_{s}$ for the standard approach (dashed lines), illustrating a performance comparison with the Wigner tomography approach (solid lines).